At what distance from this particle does its field have a magnitude of 12.0 N/C

1
Expert's answer
2022-02-22T08:53:37-0500

Solution

Let consider we have a particle of charge Q

Then electric field at distance r

"r=\sqrt{\frac{kQ}{E}}\\=\sqrt{\frac{9*10^9Q}{12}}\\=2.74\sqrt{Q}" m so put Q value we get desired solution.